OpenStudy (anonymous):

Solve the equation for a. K = 4a + 9ab

OpenStudy (dakid88):

a=K9b+4

OpenStudy (dakid88):

a=k/9b+4

OpenStudy (olyalhic):

k=4a+9ab Flip the equation. 9ab+4a=k Factor out variable a. a(9b+4)=k Divide both sides by 9b+4. \[\frac{a(9b+4) }{ 9b+4 } = \frac{ k}{ 9b+4}\] Answer \[a = \frac{ k }{ 9b+4 }\]